OpenStudy (anonymous):

7/9, 7/3, 7, -21, 63,.... 9th term of following geometric sequence? Enter a whole number

OpenStudy (campbell_st):

you need to sind the common ratio 1st \[\frac{T_{2}}{T_{1}} = \frac{T_{3}}{T_{2}}\] when you know the common ratio then \[T_{9} = ar^{n -1}\] a = 1st term 7/9 r = common ratio n = term number ( its 9)

OpenStudy (anonymous):

what do you plug in for R?

OpenStudy (campbell_st):

you need to find it... what is each term being multiplyed by to get the next term..?

OpenStudy (campbell_st):

there may be an error in to sequence, is the 2nd term negative... as you have the 4th term as a negative..

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

-3

OpenStudy (campbell_st):

7/9 * -1/3 = -7/27 so its not -1/3 just the reciprocal

OpenStudy (campbell_st):

yep thats it so \[T_{9} = \frac{7}{9} \times (-3)^{9 -1}\] the answer should be positive

OpenStudy (anonymous):

i got 878.65, is that correct? or rounded 879

OpenStudy (campbell_st):

I got 5103

OpenStudy (campbell_st):

it should be a whole number answer.